# The Prompt Makes the Person(a): A Systematic Evaluation of Sociodemographic Persona Prompting for Large Language Models

### v1

@@ E.1 Statistics on Role Violations
We analyze the extent to which a model violates its assigned role, i.e. it refuses to answer or answers with the identity of an AI assistant (see Appendix D.1 for details). We find that Llama-3.3-70B, OLMo-2-32B-Instruct, and OLMo-2-7B result in no or very few role violations (up to 2.4%). Results for Llama-3.1-8B and Gemma-3-27b are presented in Figure 7. We observe that the interview format generally leads to a higher number of role violations, with Gemma-3-27b exhibiting notable difficulties in simulating personas for certain combinations of prompt types and demographic groups. We suspect that this is due to a conflict between the persona simulation instruction and the system prompt’s directive to behave as an AI assistant.
@@ E.2 Social Media Bios
Complementary results for social media bios are presented in Figures 8, 9, and 10. The results overall align with those reported for self-descriptions.
@@ E.5 Alignment with OpinionsQA
We show the alignment of all remaining models with respect to distribution and majority match in Figure 13 and 14, respectively.
@@ F.1 Regression Analysis
Influence of Prompt Types
@@
Influence of Prompt Phrasing
To verify that our findings stem from the introduced prompt dimensions rather than specific wording, we include two alternative phrasings for each prompt template (cf. Table 3) and assess the impact of prompt phrasing on the results. We find that regression coefficients for prompt phrasing are significant in only 1 out of 5 models for the marked word count and the share of non-English responses (cf. Tables 9, 11), in 2 out 0f 5 models for distribution match (cf. Table 13), and in 3 out of 5 models for semantic diversity and accuracy (cf. Tables 10, 12). Across all measures and models, the influence (i.e., the absolute value of the regression coefficient) of using names and the interview format consistently exceeds that of the prompt phrasing, indicating that our main findings are not driven by incidental wording. We further note, that the prompt phrasing shows the least effect on Llama-3.2-70B (no significant coefficients across all measures), while OLMo2-7B is most sensitive to prompt phrasing (significant coefficients for 4 out of 5 measures).

### v2

@@ E.1 Statistics on Role Violations
We analyze the extent to which a model violates its assigned role, that is, it refuses to answer or answers with the identity of an AI assistant (see Appendix D.1 for details). We find that Llama-3.3-70B, OLMo-2-32B-Instruct, and OLMo-2-7B result in no or very few role violations (up to 2.4%). Results for Llama-3.1-8B and Gemma-3-27b are presented in Figure 7. We observe that the interview format generally leads to a higher number of role violations, with Gemma-3-27b exhibiting notable difficulties in simulating personas for certain combinations of prompt types and demographic groups. We suspect that this is due to a conflict between the persona simulation instruction and the system prompt’s directive to behave as an AI assistant.
@@ E.2 Social Media Bios
Complementary evaluation results for social media bios are presented in Figures 8, 9, and 10. We find that the results overall align with those reported for self-descriptions.
@@ E.5 Alignment with OpinionsQA
We show the opinion distance (i.e., Wasserstein distance) on OpinionsQA for all remaining models in Figure 13.
Additionally, we also report majority option match, where we compute the option with the highest probability and compute the match of this single option between models and humans. Unlike opinion distance, here higher values indicate better alignment with the human responses.171717Note that the majority option match is similar to, but not the exact same as, average accuracy over the 100 questions, since not all questions have the same answer options. We show the majority match on OpinionsQA for all models in Figure 14.
@@ F.1 Regression Analysis
Influence of Prompt Types.
@@
Influence of Prompt Phrasing.
To verify that our findings stem from the introduced prompt dimensions rather than specific wording, we include two alternative phrasings for each prompt template (cf. Table 3) and assess the impact of prompt phrasing on the results. We find that regression coefficients for prompt phrasing are significant in only 1 out of 5 models for the marked word count and the share of non-English responses (cf. Tables 9, 11), in 2 out of 5 models for opinion distance (cf. Table 13), and in 3 out of 5 models for semantic diversity and accuracy (cf. Tables 10, 12). Across all measures and models, the influence (i.e., the absolute value of the regression coefficient) of using names and the interview format consistently exceeds that of the prompt phrasing, indicating that our main findings are not driven by incidental wording. We further note, that the prompt phrasing shows the least effect on Llama-3.2-70B (no significant coefficients across all measures), while OLMo2-7B is most sensitive to prompt phrasing (significant coefficients for 4 out of 5 measures).